OpenStudy (anonymous):

If sin θ = (sqrt {3})/2, which could not be the value of θ?

OpenStudy (anonymous):

what value of \(\theta\) leads to \(\sin\theta\) being \(\underline{\sqrt3}\\2\)

OpenStudy (midhun.madhu1987):